OpenStudy (anonymous):

Samantha puts an initial $500 into a savings account. The account has a 4% annual compound interest rate. What is the function that best represents Samantha's account and how much does she have in savings after 5 years? Round your answer to the nearest dollar. Hint: Use the formula, f(x) = P(1 + r)x.

OpenStudy (anonymous):

Here are the answer options . f(x) = 500(1.4)x, $2,689 f(x) = 500(0.04)x, $512 f(x) = 500(1.04)x, $608 f(x) = 500(1.04)x, $527

OpenStudy (anonymous):

PLease help!

Vocaloid (vocaloid):

hints: r = interest rate = 0.04 P = principal amount = 500 your formula is f(x) = P(1+r)^x

OpenStudy (anonymous):

OK thank you! So would it be c?

Vocaloid (vocaloid):

not quite, we still need to check whether it's C or D when x = 5, what does 500(1.04)^x = ?

OpenStudy (anonymous):

oh ok 1 min

OpenStudy (anonymous):

It is C. Because when i put 5 for the exponent it came out to be 608.32645. Correct?

Vocaloid (vocaloid):

right! good job

OpenStudy (anonymous):

Yay! Thank you so much!